OpenStudy (anonymous):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased___adult tickets and ___child tickets.

OpenStudy (adamk):

The equations will be 9x + 4.5y = 11 and x + y = 72 where x = # of adult tickets and y = # of children tickets

OpenStudy (adamk):

Solve the systems of equations. Look at what @Nurali linked for more info.

OpenStudy (anonymous):

5 adults 6 kids

OpenStudy (legends):

she couldve purchased 8 adults and zero children because 9 times 8 equals 72

OpenStudy (anonymous):

@Legends long time no talk!

OpenStudy (adamk):

Oops my equations were wrong. 9x + 4.5y = 72 and x + y = 11.

OpenStudy (legends):

@edennoyes hey

OpenStudy (anonymous):

hi @Legends